### v2

@@ 1 Introduction
The past few years have witnessed remarkable progress in the development of Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023), giving rise to the proliferation of numerous agentic systems (Weng, 2023; Shen et al., 2024). For example, “chain-of-thought” prompting has unlocked the general-purpose reasoning capabilities of LLMs (Wei et al., 2022), and memory mechanisms have been proven effective in simulating human behavioiur (Park et al., 2023). These emerging LLM agents have demonstrated astonishing abilities to transform a wide range of tasks, including solving mathematical problems (Romera-Paredes et al., 2024), navigating the web (Nakano et al., 2021), providing financial advice (Ding et al., 2024) and informing medical decisions (Li et al., 2024a). Therefore, the design of agentic systems plays a crucial role in harnessing the power of LLMs for various downstream applications.
@@
However, current research predominantly relies on manually designed agentic systems tailored for specific tasks, which often depend heavily on expert insight and intensive human labor. Furthermore, these task-specific agent designs frequently struggle to adapt to novel tasks. A few recent works have explored using LLMs to rewrite and optimize the prompts of existing agents (Fernando et al., 2024; Yang et al., 2024). However, these rewriting approaches are limited in their ability to recombine the strengths of agentic modules discovered by different researches and located in separate codebases. Another line of research focuses on optimizing the configuration of multi-agent systems (Chen et al., 2023; Yuan et al., 2024; Li et al., 2023; Zhuge et al., ; Wang et al., 2023b). These efforts are orthogonal to the optimization of single-agent systems, as they focus more on the role-playing and interaction patterns among multiple agents, rather than the design of agentic modules.
Figure 1: AgentSquare is a modular framework for designing and optimizing LLM agents. We first propose a modular design space of LLM agents and extract 4 types of standardized modules including planning, reasoning, tooluse, and memory. Based on this, we design a novel LLM agent search framework to automatically discover good-performing agents.
@@
This paper addresses a novel research problem — Modularized LLM Agent Search (MoLAS). The goal is to automatically optimize LLM agent designs by leveraging the experience of published or evaluated modules. Therefore, the core of our work is a modular design space for LLM agents, comprising 4 categories of modules: Planning, Reasoning, Tool Use, and Memory. This design space is abstracted from a thorough literature review of existing agentic systems (details provided in Section 2). It is important to note that our goal is not to propose the most comprehensive, one-size-fits-all LLM agent design space, but rather to demonstrate that our modular design space enables researchers and intelligent search algorithms to fully exploit the potential of prior successful designs. Thus, our key consideration is providing a standardized IO interface for these modules, facilitating easy recombination of modules from different agentic systems and hence enabling efficient search for novel agents. Our design space is also highly extensible, allowing new agentic systems to be integrated as plug-in modules. Therefore, it provides a platform to consolidate the collective efforts of the research community on LLM agents. The overview of this work is illustrated in Figure 1.
@@ 2 A Modular Design Space of LLM Agents
A modular design space for LLM agents facilitates the reuse of prior successful designs and supports the exploration of new architectures. At the core of such modularization is the standardization of input-output interfaces, which ensures both extensibility and seamless integration with existing designs. Many experts in the field have proposed building LLM agentic systems with key modular components from engineering (Weng, 2023) and cognitive perspectives (Sumers et al., 2023; Shao et al., 2024). However, these proposals remain largely conceptual, lacking implementable solutions to unify existing LLM agents. Besides, current LLM workflow program frameworks (e.g., LangChain and AutoGPT) only provide operation-level components, which cannot support module-level search that best exploits the potential of prior successful designs.
@@
To address these problems, we perform a comprehensive literature review of publications from NeurIPS, ICML, and ICLR over the past three years. The review focuses on papers with the keywords “LLM”, “Agent”, or “Large Language Model” in their titles while excluding works related to multi-agent systems or agents that require additional training. Note that our aim is not to propose the most comprehensive, one-for-all LLM agent design space, but to offer a standardized framework that enables the recombination of existing agents and facilitates the discovery of new ones. As a result, we sort out 16 popular LLM agents and abstract a modular design space with 1050 possible combinations, which can be easily extended when new modules are discovered. Below, we describe the agentic workflow and the function of each modules in our design space.
@@
Workflow overview. The proposed agent workflow operates through an iterative process with the interconnection of the above four modules, as shown in Figure 2. Upon receiving a task $d$, the agent starts with the planning module, decomposing it into $n$ sub-tasks$\{s_{1},s_{2},\dots,s_{n}\}$. Next, these sub-tasks are passed to the reasoning module sequentially. Taking the sub-task $s_{i}$ description as input, the reasoning module explores to prompt LLMs to give the result. When reasoning encounters limitations in internal knowledge of LLMs, the tool use module is activated to select an appropriate tool from the pre-defined tool pool $\tau$, supporting problem-solving. Besides, the reasoning process also accesses the memory module which reads and writes necessary observations and experiences from a memory database $mem$ to help reasoning. The reasoning result of each sub-task will be transformed into actions, guiding the agent to interact with the external environment. After all sub-tasks are finished or the reasoning process gets stacked, the agent will activate the planning module to adjust the plan with the received feedback. The agent conducts such a trial-and-error loop until the task $d$ is completed or the set maximum trial number is reached.
@@ 3.2 AgentSquare search algorithm
Output: The evolved agent $A^{*}$
$t\leftarrow 1$ // Current search episode
$A_{e}^{0}\leftarrow A_{0}$ // Initialization of the module evolution phase
while $t\leq K$ do
       $t\leftarrow t+1$
$A^{*}\leftarrow A_{e}^{0}$
return $A^{*}$
Algorithm 1 Algorithm of AgentSquare
@@ 3.5 Module Evolution
As mentioned above, the solution space for each module type is open-ended, allowing any code with a standardized I/O interface. Consequently, searching only with module recombination narrows the solution space and limits the upper bound of agent performance. To address this problem, we design a module evolution operation with an evolutionary meta-prompt to search for new modules through program-level optimization. Specifically, we introduce a module-programming LLM $\pi_{\xi}$ to evolve new modules by jointly modeling task descriptions, existing modules, and the performance of previously evaluated modules. Leveraging LLMs as module programmer has several appealing advantages. Modern LLMs are typically trained on extensive code datasets, enabling advanced code comprehension and generation capabilities. Additionally, integrating existing module codes as in-context examples, along with past experiences, can elicit the reflective reasoning abilities of LLMs to identify promising optimization directions. Denote the initial agent in the module evolution stage as $A_{e}^{0}=(P^{{}^{\prime}}_{0},R^{{}^{\prime}}_{0},T^{{}^{\prime}}_{0},M^{{}^{\prime}}_{0})$, the module programmer LLM produces a population of child agents by evolving current modules of $A_{e}^{0}$. Formally the module evolution operation is denoted as follows:
@@ 4.1 Experimental Setup
Task setup. We conduct experiments on six representative tasks covering four domains: embodied, game, web and tool applications, which are widely adopted by existing LLM agent benchmarks (Ma et al., 2024a; Xi et al., 2024), more details are presented in Appendix A.1.
Baselines. We compare AgentSquare with three types of baselines including hand-crafted agents, module-level search, and prompt-level search methods. More details are presented in Appendix A.1.
@@
Baseline Type Method Webshop ALFWorld SciWorld M3Tool TravelPlanner PDDL
@@
Table 1: Performance comparison of searched agents from AgentSquare and (1) existing human-designed agents (2) module search baselines (3) prompt search baselines based on GPT-4o on six tasks across different domains.
@@ 4.2 Experimental Results
  • AgentSquare provides a more efficient searching approach for LLM agent optimization. To further demonstrate the effectiveness of the search pipeline of AgentSquare, we also compare with two types of agent searching methods including module-level and prompt-level search. Compared with the best agent crafted from these searching methods, AgentSquare achieves an average 8.4% performance improvement on Webshop, 8.1% improvement on ALFWorld, 11.0% improvement on SciWorld, 12.8% improvement on M3Tool, 2.5% improvement on Travelplanner, 1.4% improvement on PDDL.
@@ 4.3 Ablation Study of AgentSquare
Effectiveness of module evolution and recombination. There are two key operations in the searching framework of AgentSquare: module evolution which creates new modules and module recombination which strategically recombines existing ones. To verify the effectiveness of each design, we tested three variants of AgentSquare: the full model, a version without module evolution, and a version without module recombination. The results based on GPT-4o and GPT-3.5 are presented in Table 2 and Table A.5, respectively. It can be seen that dropping each design results in a noticeable performance decline and the module recombination has a larger impact. Module recombination significantly expands the search space, reducing the risk of falling into a local optima. Meanwhile, module evolution facilitates the discovery of more advanced modules tailored to specific tasks. These two operations collaborate well ensuring the effectiveness of the search process in AgentSquare.
@@
Effectiveness of performance predictor. In this part, we verify the effectiveness of this design empirically. Figure 5 illustrates the predicted performance of given agents versus their actual tested performance based on both GPT-3.5 and GPT-4o on all six tasks. It can be found that the predicted performance closely aligns with the actual performance, demonstrating the effectiveness of the performance predictor. For instance, the evaluation cost of the predictor is only about 0.025% of the cost of a full evaluation based on GPT-4o in ALFWorld, demonstrating its remarkable cost-efficiency.
@@ 5.1 LLM-based Autonomous Agents
LLM-based autonomous agents are an advanced AI system using a core LLM to manage external functional modules and interact with environments. In the early stage, LLM-based agents possess very simple architectures, which just directly query LLMs to conduct decision-making and take action (Huang et al., 2022; Ahn et al., 2022). More recently, advancements have equipped agents with some other LLM-centric functional modules including planning (Hao et al., 2023; Huang et al., 2023; Zeng et al., 2024), reasoning (Wei et al., 2022; Yao et al., 2024; Shang et al., 2024), using tools (Shen et al., 2024; Schick et al., 2024), and monitoring memory (Wang et al., 2024a; Park et al., 2023), greatly enhancing the capabilities of LLM agents. Along with the improvement of the single agent, there’s another line of work trying to build more advanced multi-agent systems by strategically organizing individual agents for both simulation (Li et al., 2023; Chen et al., 2023) and targeted task solving (Qian et al., 2024; Xiao et al., 2023; Chen et al., 2024b; Li et al., 2024b). The emergence of more and more sophisticated agent produces remarkable performance improvement, however, their architectures and codebases differ greatly with each other. The lack of a unified design space and consistent terminologies across individual works makes it hard to compare different agents, understand their evolution routes, and guide new agent design directions.
@@ 5.2 Automatic Design of LLM-based Agents
LLM-based agent system, as the most advanced AI system, has not yet formed a unified design space and an automatic design approach. Existing efforts on automatic building of LLM agents focus on three aspects: engineering-oriented open resources, conceptual framework of LLM agent design, and automatic design of partial components of LLM agents. Firstly, engineering-oriented open resources like LangChain***https://github.com/langchain-ai/langchain and BabyAGIhttps://github.com/yoheinakajima/babyagi has provided convenient ways to build an LLM-centric agentic system, however, they still need human participation to organize different modules and can’t support the optimization of the designed agent. Secondly, there’s has been some conceptual framework trying to provide a unified design principle of LLM agents, such as CoALA (Sumers et al., 2023). It points out four basic modules of language agents from the perspective of cognitive science, including long-term memory, external grounding, internal actions and decision-making. However, it’s still a vision of how LLM agents should be in the future, without providing a practical design framework. Thirdly, there have been some separate works to automate parts of LLM agent systems, such as prompt optimization (Do et al., 2024; Ma et al., 2024b) and workflow search (Li et al., 2024c; Hu et al., 2024). Nevertheless, these works over-simplify the structure of agentic systems to pure prompts or codes, unable to provide a clear design space of LLM agents. Besides, another common shortcoming of existing works is that all of them are proposed without much consideration of existing designed agent architectures, thus hard to be representative and universal. Differently, our proposed modularized LLM agent design space can well cover existing efforts of LLM agent design, and provide a practical approach for automatic agent architecture optimization.
@@ 6 Conclusion
In this work, we introduce a novel modular design space for LLM agents, allowing researchers to build upon successful prior designs and collectively accumulate new insights. Based on this, we propose a novel research problem, Modularized LLM Agent Search (MoLAS), which aims to automatically optimize LLM agent designs by leveraging the knowledge gained from previously published or evaluated modules. To address the challenge of vast search spaces, we present AgentSquare, an automatic search framework to optimize LLM agents through module evolution and recombination, enabling the exploration of the entire solution space. We further introduce a performance predictor as an in-context surrogate model for evaluating novel LLM agents to accelerate the search process. AgentSquare consistently discovers novel agents that surpass all existing hand-crafted ones, while also offering interpretable design insights that deepen our understanding of these new agents. Overall, our work offers a transition from studying individual LLM agent designs to studying LLM agents within a modular design space, further consolidating the collective efforts of research community.

### v3

@@ 1 Introduction
The past few years have witnessed remarkable progress in the development of Large Language Models (LLMs) (Achiam et al., 2023; Touvron et al., 2023), giving rise to the proliferation of numerous agentic systems (Weng, 2023; Shen et al., 2024). For example, “chain-of-thought” prompting has unlocked the general-purpose reasoning capabilities of LLMs (Wei et al., 2022), and memory mechanisms have been proven effective in simulating human behavioiur (Park et al., 2023). These emerging LLM agents have demonstrated astonishing abilities to transform a wide range of tasks, including solving mathematical problems (Romera-Paredes et al., 2024), navigating the web (Nakano et al., 2021), providing financial advice (Ding et al., 2024a) and informing medical decisions (Li et al., 2024a). Therefore, the design of agentic systems plays a crucial role in harnessing the power of LLMs for various downstream applications.
@@
However, current research predominantly relies on manually designed agentic systems tailored for specific tasks, which often depend heavily on expert insight and intensive human labor. Furthermore, these task-specific agent designs frequently struggle to adapt to novel tasks. A few recent studies have explored using LLMs to rewrite and optimize the prompts of existing agents (Fernando et al., 2024; Yang et al., 2024). A more recent work introduces the idea to leverage LLMs to search the entire agentic systems defined in code space (Hu et al., 2024), enabling the discovery of agents with more flexible prompts, control flows, etc. However, these previous approaches are limited in their ability to explicitly recombine the strengths of agentic modules discovered by different researches and located in separate codebases. Another line of research focuses on optimizing the configuration of multi-agent systems (Chen et al., 2023; Yuan et al., 2024; Li et al., 2023; Zhuge et al., 2024; Wang et al., 2023b). These efforts are orthogonal to the optimization of single-agent systems, as they focus more on the role-playing and interaction patterns among multiple agents, rather than the design of agentic modules.
Figure 1: AgentSquare is a modular framework for designing and optimizing LLM agents.
@@
This paper addresses a novel research problem — Modularized LLM Agent Search (MoLAS). The goal is to automatically optimize LLM agent designs by leveraging the experience of published or evaluated modules. Therefore, the core of our work is a modular design space for LLM agents, comprising 4 categories of modules: Planning, Reasoning, Tool Use, and Memory. This design space is abstracted from a thorough literature review of existing agentic systems (details provided in Section 2). It is important to note that our goal is not to propose the most comprehensive, one-size-fits-all LLM agent design space, but rather to demonstrate that our modular design space enables researchers and intelligent search algorithms to fully exploit the potential of prior successful designs. MoLAS is a guided and constrained searching problem in the modular design space, which is a subset of the entire code search proposed in ADAS (Hu et al., 2024). However, MoLAS has a nice feature of providing standardized IO interfaces for agent modules, facilitating easy recombination of modules from different agentic systems and hence enabling efficient search for novel agents. Our design space is also highly extensible, allowing new agentic systems to be integrated as plug-in modules. Therefore, it provides a platform to consolidate the collective efforts of the research community on LLM agents. The overview of this work is illustrated in Figure 1.
@@ 2 A Modular Design Space of LLM Agents
2.1 Background
Using LLMs for automatic optimization has been a widely explored topic, such as applications in code generation (Lehman et al., 2023; Romera-Paredes et al., 2024) and neural architecture search (Nasir et al., 2024; Chen et al., 2024a). There are several recent studies that explore the problem of prompting LLMs to design LLM agentic systems. OPRO (Yang et al., 2024) and Promptbreeder (Fernando et al., 2024) can be viewed as leveraging the reasoning power of LLMs to improve the prompt of LLM agents. More importantly, ADAS introduces the idea of searching the entire agentic system defined in code space, and propose a Meta Agent Search algorithm that discovers LLM agents outperforming state-of-the-art human designs (Hu et al., 2024). Our main difference and contribution lie in introducing a modular design space for LLM agents, which can provide a standard framework to support the convenient reuse of existing successful agent components and fruitful innovative agent module discovery.
@@ 2.1 Background
A modular design space for LLM agents facilitates the reuse of prior successful designs and supports the exploration of new architectures. At the core of such modularization is the standardization of input-output interfaces, which ensures both extensibility and seamless integration with existing designs. Many experts in the field have proposed building LLM agentic systems with key modular components from engineering (Weng, 2023) and cognitive perspectives (Sumers et al., 2023). However, these proposals remain largely conceptual, lacking implementable solutions to unify existing LLM agents. Besides, current LLM workflow program frameworks (e.g., LangChain and AutoGPT) only provide operation-level components, which cannot support module-level search that best exploits the potential of prior successful designs.
@@
To address these problems, we perform a comprehensive literature review of publications from NeurIPS, ICML, and ICLR over the past three years. The review focuses on papers with the keywords “LLM”, “Agent”, or “Large Language Model” in their titles while excluding works related to multi-agent systems or agents that require additional training. Note that our aim is not to propose the most comprehensive, one-for-all LLM agent design space, but to offer a standardized framework that enables the recombination of existing agents and facilitates the discovery of new ones. As a result, we sort out 16 popular LLM agents and abstract a modular design space with 1050 possible combinations, which can be easily extended when new modules are discovered. Below, we describe the agentic workflow and the function of four modules in our design space.
2.2 Workflow overview
@@
The proposed agent workflow operates through an iterative process with the interconnection of the above four modules, as shown in Figure 2. Upon receiving a task $d$, the agent starts with the planning module, decomposing it into $n$ sub-tasks$\{s_{1},s_{2},\dots,s_{n}\}$. Next, these sub-tasks are passed to the reasoning module sequentially. Taking the sub-task $s_{i}$ description as input, the reasoning module explores to prompt LLMs to give the result. When reasoning encounters limitations in internal knowledge of LLMs, the tool use module is activated to select an appropriate tool from the pre-defined tool pool $\tau$, supporting problem-solving. Besides, the reasoning process also accesses the memory module which reads and writes necessary observations and experiences from a memory database $mem$ to help reasoning. The reasoning result of each sub-task will be transformed into actions, guiding the agent to interact with the external environment. After all sub-tasks are finished or the reasoning process gets stacked, the agent will activate the planning module to adjust the plan with the received feedback. The agent conducts such a trial-and-error loop until the task $d$ is completed or the set maximum trial number is reached.
@@ 3.5 Module Evolution
As mentioned above, the solution space for each module type is open-ended, allowing any code with a standardized I/O interface. Consequently, searching only with module recombination narrows the solution space and limits the upper bound of agent performance. To address this problem, we design a module evolution operation with an evolutionary meta-prompt to search for new modules through program-level optimization. This design is inspired by the iterative pipeline of FunSearch (Romera-Paredes et al., 2024), which prompts LLMs to propose new solutions based on the target problem and performance feedback from existing solutions. Building on this concept, we introduce a module-programming LLM $\pi_{\xi}$ to conduct agent search in our modular design space by jointly modeling task descriptions, existing modules, and the performance of previously evaluated modules. Please note we reuse parts of the open-source code from ADAS (Hu et al., 2024) to implement the optimization procedure. Leveraging LLMs to search in the modular agent design space has several appealing advantages. Compared with the unconstrained design space of LLM agents, searching functional modules can produce a more focused and fruitful search space. Additionally, integrating existing successful module designs with standard IO as in-context examples can better elicit the reflective reasoning abilities of LLMs to identify previous key designs to help propose innovative ones. Denote the initial agent in the module evolution stage as $A_{e}^{0}=(P^{{}^{\prime}}_{0},R^{{}^{\prime}}_{0},T^{{}^{\prime}}_{0},M^{{}^{\prime}}_{0})$, the module programmer LLM produces a population of child agents by evolving current modules of $A_{e}^{0}$. Formally the module evolution operation is denoted as follows:
@@ 4.1 Experimental Setup
Task setup. We conduct experiments on six representative tasks covering four domains: embodied, game, web and tool applications, which are widely adopted by existing LLM agent benchmarks (Ma et al., 2024; Xi et al., 2024), more details are presented in Appendix A.1.
Baselines. We compare AgentSquare with four types of baselines including hand-crafted agents, module-level search, prompt-level search and agent-search methods. More details are presented in Appendix A.1.
@@
Baseline Type Method Webshop ALFWorld SciWorld M3Tool Travel PDDL
@@
Table 1: Performance comparison of searched agents from AgentSquare and (1) existing human-designed agents (2) module search baselines (3) prompt search baselines (4) agent search baselines based on GPT-4o on six tasks across different domains.
@@ 4.2 Experimental Results
  • AgentSquare provides a more efficient searching approach for LLM agent optimization. To further demonstrate the effectiveness of the search of AgentSquare, we compare three types of searching methods including module search, prompt search and agent search. Compared with the best agent crafted from these searching methods, AgentSquare achieves an average 8.4% performance improvement on Webshop, 8.1% improvement on ALFWorld, 11.0% improvement on SciWorld, 12.8% improvement on M3Tool, 2.5% improvement on Travelplanner, 1.4% improvement on PDDL. The comparison of search-based methods is conducted with a fixed LLM token budget to ensure fairness by maintaining the same number of search iterations. While in principle ADAS has the potential to discover more sophisticated agents by searching in the entire code space, it may require more iterations (and thus higher LLM token usage) to achieve this.
@@ 4.3 Ablation Study of AgentSquare
Effectiveness of module evolution and recombination. There are two key operations in the searching framework of AgentSquare: module evolution which creates new modules and module recombination which strategically recombines existing ones. To verify the effectiveness of each design, we tested three variants: the full model, a version without module evolution, and a version without module recombination. The results based on GPT-4o and GPT-3.5 are presented in Table 2 and Table A.5, respectively. It can be seen that dropping each design results in a noticeable performance decline and the module recombination has a larger impact. Module recombination significantly expands the search space, reducing the risk of falling into a local optima. Meanwhile, module evolution facilitates the discovery of more advanced modules tailored to specific tasks. These two operations collaborate well ensuring the effectiveness of the search process in AgentSquare.
@@
Effectiveness of performance predictor. In this part, we verify the effectiveness of this design empirically. Figure 5 illustrates the predicted performance of given agents versus their actual tested performance based on both GPT-3.5 and GPT-4o on all six tasks. The tested agents were generated through random sampling by randomly combining existing modules. It can be found that the predicted performance closely aligns with the actual performance, demonstrating the effectiveness of the performance predictor. For instance, the evaluation cost of the predictor is only about 0.025% of the cost of a full evaluation based on GPT-4o in ALFWorld, demonstrating its remarkable cost-efficiency. We provide more experiment results of predicting the performance of the dynamically searched agents in Figure A.15 of the Appendix.
@@ 5.1 LLM-based Autonomous Agents
LLM-based autonomous agents are an advanced AI system using a core LLM to manage external functional modules and interact with the world (Ding et al., 2024b). Recent studies have equipped LLM agents with several LLM-centric functional modules including planning (Hao et al., 2023; Zeng et al., 2024; Shao et al., 2025), reasoning (Wei et al., 2022; Yao et al., 2024; Shang et al., 2024; Xu et al., 2025), using tools (Shen et al., 2024; Schick et al., 2024), and monitoring memory (Wang et al., 2024a; Park et al., 2023), greatly enhancing the capabilities of LLM agents. Along with the improvement of the single agent, there’s another line of work trying to build more advanced multi-agent systems by strategically organizing individual agents for both simulation (Li et al., 2023; Chen et al., 2023) and targeted task solving (Qian et al., 2024; Chen et al., 2024b; Li et al., 2024b). The emergence of more and more sophisticated agent produces remarkable performance improvement, however, their architectures and codebases differ greatly with each other. The lack of a unified design space and consistent terminologies across individual works makes it hard to compare different agents, understand their evolution routes, and guide new agent design directions.
@@ 5.2 Automatic Design of LLM-based Agents
LLM-based agent system, as the most advanced AI system, has not yet formed a unified design space and an automatic design approach. Engineering-oriented open resources like LangChain***https://github.com/langchain-ai/langchain and BabyAGIhttps://github.com/yoheinakajima/babyagi have provided convenient ways to build an LLM-centric agentic system, however, they still need human participation to organize different modules and can’t support the optimization of the designed agent. Besides, there have been some conceptual frameworks trying to provide a unified design principle of LLM agents, such as CoALA (Sumers et al., 2023). However, it’s still a vision of how LLM agents should be in the future, without providing a practical design framework. More importantly, there are several recent works that explore the problem of automating (at least part of) the design of LLM agent systems defined on different search spaces. OPRO (Yang et al., 2024) and Promptbreeder (Fernando et al., 2024) can be considered as using LLMs to optimize LLM agent defined on prompt space. More relevantly, ADAS (Hu et al., 2024) proposes to search the entire agentic system defined on code space, enabling the search for LLM agents with more flexible prompts, tool uses, control flows and more.
@@ 6 Conclusion
In this work, we introduce a novel modular design space for LLM agents, allowing researchers to build upon successful prior designs and collectively accumulate new insights. Based on this, we propose a novel research problem, Modularized LLM Agent Search (MoLAS), which aims to automatically optimize LLM agent designs by leveraging the knowledge gained from previously published or evaluated modules. To address the challenge of vast search spaces, we present AgentSquare, an automatic search framework to optimize LLM agents through module evolution and recombination. We further introduce a performance predictor as an in-context surrogate model for evaluating novel LLM agents to accelerate the search process. Overall, our work offers a transition from studying individual LLM agent designs to studying LLM agents within a modular design space, further consolidating the collective efforts of the research community.